OpenStudy (anonymous):

Perform the requested operation or operations. f(x) = (x-3)/(8) .; g(x) = 8x + 3, find g(f(x)).

OpenStudy (anonymous):

Answer Choices: g(f(x)) = x g(f(x)) = x - three divided by eight. g(f(x)) = 8x + 21 g(f(x)) = x + 6

OpenStudy (anonymous):

g(f(x)) = x

OpenStudy (anonymous):

The Answer is \(g(f(x))=x\) \[g(x)=8x+3\] \[g(f(x))=8\left(\frac{x-3}{8}\right)+3=(x-3)+3=x-3+3=x\]

OpenStudy (anonymous):

See, that is the answer I got. Nonetheless, I just was not sure. However, thank you for actually demonstrating how to solve the problem @KeithAfasCalcLover . I am very grateful.

OpenStudy (anonymous):

@Omistone Thank You as well :D

OpenStudy (anonymous):

My Pleasure

OpenStudy (anonymous):

:)